# Relaxation of first-class constraints and the quantization of gauge theories: from “matter without matter” to the reappearance of time in quantum gravity

### v1

@@ II.2 Gauge fixing
$\phi_{a}(z)$ $\approx 0\ ,\ \chi^{a}(\tau;z,\lambda]\approx 0\ ,$
@@
A useful particular case is that in which the gauge-fixing conditions are only functions of the $z$ fields, $\chi^{a}(\tau;z,\lambda]\equiv\chi^{a}(\tau;z(\tau))$. These are the so-called ‘canonical gauge-fixing conditions’ or simply ‘canonical gauges’ Hanson ; HT:book . Then, taking $p=2$ without loss of generality, the conservation equation
@@
$\lambda^{a}\approx-\left(\Delta^{-1}\right)^{a}_{b}\left(\frac{\partial\chi^{b}}{\partial\tau}+\{\chi^{b},H_{0}\}\right)\equiv\Lambda^{a}(\tau;z)\ ,$ (25)
@@ III.1.2 Noncanonical gauges and the relaxation of the parent theory’s constraints
which is simply obtained from the action (4) by imposing $\lambda^{a_{u}}=\Lambda^{a_{u}}(\tau;z]$ before variation. From the point of view of the parent theory, this is an erroneous gauge-fixing procedure, whereas it is a mechanism to relax the constraints and introduce new degrees of freedom in the Fock–Stueckelberg theory. [^13] As in the parent gauge theory, the Rosenfeld–Bergmann–Dirac algorithm Rosenfeld ; SalisburyRosenfeld ; Dirac3 ; RBD ; HT:book ; Bergmann that ensures the theory’s consistency by verifying that the constraints hold at all times should also be applied to the Fock–Stueckelberg theory. This means that we must require that the time derivatives of the constraints derived, for example, from the action given in Eq. (31) vanish on shell. This may lead to a fixation of multipliers $\lambda$, to redudant constraints or to new constraints. In case new constraints appear, we must again require they hold at all times.
@@ III.2 Constants of Nature as spacetime fields
$S_{\rm new}$ $=\int_{\tau_{0}}^{\tau_{1}}\!\mathrm{d}\tau\Big{\{}\mathcal{A}_{\mu}(z,\mathscr{P})\dot{z}^{\mu}+\mathscr{P}_{i}\dot{\mathscr{Q}}^{i}$ (32)
$\ \ \ -H_{0}(\tau;z,\mathscr{P}]-\lambda^{a}\phi_{a}(z,\mathscr{P})\Big{\}}\,,$
which is identical to Eq. (4) except for the fact that the constants of Nature $\mathscr{P}_{i}$ are promoted to fields (degrees of freedom), which are interpreted as the canonical momenta conjugate to cyclic degrees of freedom $\mathscr{Q}^{i}$. This then implies that the $\mathscr{P}$ fields do not change over time, and any initial condition will be frozen. If one assumes an initial value for the $\mathscr{P}$ fields that is constant across space (if $d>0$), then due to $\mathscr{P}$ conservation, the $\mathscr{P}$ solution to the field equations will be spatially constant at all times. On the other hand, an initial condition that changes across space would lead to “constants” of Nature with a static spatial variation. In any case, the phase space is extended to include the canonical pairs $(\mathscr{Q},\mathscr{P})$, and the new theory described by the action (32) does not exhibit constants of Nature (theory-defining parameters) but rather only fields, some of which are conserved in time. Of course, we also assume that the constraints $\phi_{a}$ obey the regularity condition with respect to the enlarged set of fields $(z,\mathscr{P},\mathscr{Q})$.
@@ III.2.1 Parametrization
First, if $d>0$, we assume that the unconstrained Hamiltonian can be written as (cf. Footnote 2)
$H_{0}(\tau;z,\mathscr{P}]=\int\mathrm{d}^{d}x\ \mathscr{H}_{0}(\tau,\mathbf{x};z(\tau,\mathbf{x}),\mathscr{P})\ ,$ (33)
@@
$\{\phi_{a}(z(\tau,\mathbf{x}),\mathscr{P}),\mathscr{H}_{0}(\tau,\mathbf{x}^{\prime};z(\tau,\mathbf{x}^{\prime}),\mathscr{P})\}$ (34)
$\ \ =\sum_{b=1}^{N}\int\mathrm{d}^{d}y\,\mathscr{V}_{a}^{b}(\tau,\mathbf{x},\mathbf{x}^{\prime},\mathbf{y};z,\mathscr{P}]\phi_{b}(z(\tau,\mathbf{y}),\mathscr{P})\ ,$
@@
then Eqs. (33) and (34) yield
$\{\phi_{a}(z(\tau,\mathbf{x}),\mathscr{P}),H_{0}(\tau;z,\mathscr{P}]\}$ (36)
$\ \ \ =\sum_{b=1}^{N}\int\mathrm{d}^{d}y\,\mathcal{V}_{a}^{b}(\tau,\mathbf{x},\mathbf{y};z,\mathscr{P}]\phi_{b}(z(\tau,\mathbf{y}),\mathscr{P})\ ,$
@@
$\phi_{0}=p_{T}+\mathscr{H}_{0}(T;\mathbf{x};z,\mathscr{P})\ ,$ (37)
@@
If we include the $(T,p_{T})$ pair in the definition of the Poisson bracket in the usual way, it is straightforward to see that $\phi_{0}$ is a first-class constraint due to the original first-class algebra given by Eqs. (12) and (34), and due to the fact that the $\phi_{a}$ constraints are assumed to have no explicit dependence on time (or $T$). We can then redefine the multipliers $\lambda_{\rm new}\coloneqq(\lambda^{0},\lambda_{\rm old})$ and constraints $\phi_{\rm new}=(\phi_{0},\phi_{\rm old})$, so as to obtain a new first-class algebra with no unconstrained Hamiltonian. As a matter of notation, we can also add $p_{T}$ to the list of $\mathcal{P}$ fields, and, correspondingly $T$ to the list of $\mathcal{Q}$ fields. [^15] This is just for convenience because, in general, $T$ will not be cyclic and $p_{T}$ will not be constant in time. If $\mathscr{H}_{0}$ or $H_{0}$ depend on $T$ explicitly, we have $\dot{p}_{T}(\tau,\mathbf{x})\approx-\int\mathrm{d}^{d}y\lambda^{0}(\tau,\mathbf{y})\delta\mathscr{H}_{0}|_{\mathbf{y}}/\delta T(\tau,\mathbf{x})$ if $d>0$ or $\dot{p}_{T}(\tau)\approx-\lambda^{0}(\tau)\partial H_{0}/\partial T$ if $d=0$. In this way, the parametrized action in Eq. (39) is equivalent to
$S_{\rm new}^{\rm par}$ $=\int_{\tau_{0}}^{\tau_{1}}\!\mathrm{d}\tau\left[\mathcal{A}_{\mu}(z)\dot{z}^{\mu}+\mathscr{P}_{i}\dot{\mathscr{Q}}^{i}-\lambda^{a}\phi_{a}(z,\mathscr{P})\right],$ (40)
where the indices now run over the appropriate values for the enlarged set of variables together with implicit spatial integrations if $d>0$. We also assume that the enlarged set of constraints obeys the regularity condition with respect to the $(z,\mathscr{P},\mathscr{Q})$ fields.
@@ III.2.2 Solving the constraints in terms of constants of Nature
$\mathscr{P}_{a}\approx-\mathscr{H}_{a}(z,\mathscr{P}_{i>a})\ ,\ (a=1,\ldots,P)\ ,$ (41)
in a region of phase space, where the constraint hypersurface $\Sigma$ can be defined by the new set of constraints [^18] The parametrization constraint in Eq. (37) is not necessarily a particular case of Eq. (42) because $p_{T}$ need not be constant in time, i.e., a constant of Nature (cf. Footnote 15). Furthermore, notice that while the constraints that come from the parametrization procedure, such as the one in Eq. (37), may be globally valid in phase space, the solutions in Eqs. (42) may hold only in a certain region. Thus, in general, a parametrized theory with constraints of the form given in Eq. (42) may not be globally equivalent to gauge theory with constraints $\phi_{a}$, with the equivalence holding locally in certain cases. We also disregard a possible explicit spatial dependence of $\mathscr{H}_{a}$ in Eqs. (41) and (42).
$\bar{\phi}_{a}$ $\coloneqq\mathscr{P}_{a}+\mathscr{H}_{a}(z,\mathscr{P}_{i>a})\approx 0\ ,\ (a=1,\ldots,P)\ ,$ (42)
@@ III.2.3 General constraint redefinitions
What can we conclude about the constraint transformation in Eqs. (42)? If we assume that this transformation is invertible in a region of phase space in a neighborhood of $\Sigma$, then it is a particular case of a general transformation $\phi_{a}\mapsto\bar{\phi}_{a}(\phi)$ of constraints that is well defined and invertible in a neighborhood of $\Sigma$ and satisfies $\bar{\phi}_{a}=0$ if, and only if, $\phi_{a}=0$ (for every $a=1,\ldots,N$). Any such transformation must satisfy $\bar{\phi}_{a}=\tensor{M}{{}_{a}^{b}}\phi_{b}$ due to Eq. (11), with $\tensor{M}{{}_{a}^{b}}$ being the coefficients of a matrix that must be non-divergent and invertible at least in a neighborhood of $\Sigma$.
The new constraints $\bar{\phi}_{a}$ are thus independent, and, by denoting all the fields by $\xi^{\mu}=(z^{\nu},\mathscr{Q}^{i},\mathscr{P}_{j})$, we obtain the gradient $\partial\bar{\phi}_{a}/\partial\xi^{\mu}\approx\tensor{M}{{}_{a}^{b}}\partial\phi_{b}/\partial\xi^{\mu}$, which must have the same rank as the matrix given by $\partial\phi_{b}/\partial\xi^{\mu}$ because $\tensor{M}{{}_{a}^{b}}$ corresponds to a full-rank matrix. In this way, the regularity condition of the constraints $\phi_{a}$ is preserved. The new constraints $\bar{\phi}_{a}\approx 0$ determine the same hypersurface $\Sigma$ as the original $\phi_{a}$ functions, but they will generally have different structure functions.
@@
As a particular case, we assume that the new constraints in Eqs. (42) are functionally independent, and that they obey the regularity condition because the gradients of $\bar{\phi}_{a}$ on $\Sigma$ are all linearly independent where Eq. (41) is valid. To verify the relation $\bar{\phi}_{a}=\tensor{M}{{}_{a}^{b}}\phi_{b}$, one may expand $\phi_{a}(z,\mathscr{P})$ ($a=1,\ldots,P$) in a Taylor series in powers of $\mathscr{P}$ to find, in particular,
$\phi_{a}=\phi_{a}(-\mathscr{H})+\frac{\partial\phi_{a}}{\partial\mathscr{P}_{b}}\bar{\phi}_{b}+\frac{1}{2}\frac{\partial^{2}\phi_{a}}{\partial\mathscr{P}_{b}\mathscr{P}_{c}}\bar{\phi}_{b}\bar{\phi}_{c}+\mathcal{O}(\bar{\phi}^{3})\,,$ (43)
where $a,b=1,\ldots,P$, $\bar{\phi}_{a}$ are given by Eq. (42), and we suppressed the arguments of $\phi_{a}$ other than $\mathscr{P}$ for brevity. Then, since $\phi_{a}(-\mathscr{H})=0$, we have
@@
As we assume that $\phi_{a}\approx 0$ ($a=1,\ldots,P$) if, and only if, the constraints in Eq. (41) hold, then the matrix $\tensor{\left(M^{-1}\right)}{{}_{a}^{b}}$ must be invertible on $\Sigma$. In a neighborhood of $\Sigma$, for $a,b,c=1,\ldots,P$, we can define $\frac{1}{2}\partial^{2}\phi_{a}/\partial\mathscr{P}_{b}\partial\mathscr{P}_{c}$ to be the element of row $a$ and column $b$ of a matrix $\mathtt{H}^{c}$ for each value of $c$, so that $\det M^{-1}=\det M^{-1}|_{\bar{\phi}=0}(1+\bar{\phi}_{c}\mathrm{Tr}(M\mathtt{H}^{c})|_{\bar{\phi}=0}+\mathcal{O}(\bar{\phi}^{2}))$, which is nonvanishing for $\bar{\phi}$ with sufficiently small but nonzero absolute values. We can then trivially extend $\tensor{(M^{-1})}{{}_{a}^{b}}$ to $a,b=1,\ldots,N$ because the constraints $\phi_{a}$ with $a=P+1,\ldots,N$ are not changed in this particular case. The theory can thus be defined from the action
$S_{\rm new}^{\rm depar}$ $=\int_{\tau_{0}}^{\tau_{1}}\!\mathrm{d}\tau\left[\mathcal{A}_{\mu}(z)\dot{z}^{\mu}+\mathscr{P}_{i}\dot{\mathscr{Q}}^{i}-\bar{\lambda}^{a}\bar{\phi}_{a}(z,\mathscr{P})\right],$ (45)

### v2

@@ II.2 Gauge fixing
$\phi_{a}$ $\approx 0\ ,\ \chi^{a}(\tau;z,\lambda]\approx 0\ ,$
@@
A useful particular case is that in which the gauge-fixing conditions only depend on the $z$ fields, $\chi^{a}(\tau;z,\lambda]\equiv\chi^{a}(\tau;z]$. These are called ‘canonical gauge-fixing conditions’ or simply ‘canonical gauges’ Hanson ; HT:book . Then, taking $p=2$ without loss of generality, the conservation equation
@@
$\lambda^{a}\approx-\left(\Delta^{-1}\right)^{a}_{b}\left(\frac{\partial\chi^{b}}{\partial\tau}+\{\chi^{b},H_{0}\}\right)\equiv\Lambda^{a}(\tau;z]\ ,$ (25)
@@ III.1.2 Noncanonical gauges and the relaxation of the parent theory’s constraints
which is simply obtained from the action (4) by imposing $\lambda^{a_{u}}=\Lambda^{a_{u}}(\tau;z]$ before variation. From the point of view of the parent theory, this is an erroneous gauge-fixing procedure, whereas it is a mechanism to relax the constraints and introduce new degrees of freedom in the Fock–Stueckelberg theory. [^13] As in the parent gauge theory, the Rosenfeld–Bergmann–Dirac algorithm Rosenfeld ; SalisburyRosenfeld ; Dirac3 ; RBD ; HT:book ; Bergmann that ensures the theory’s consistency by verifying that the constraints hold at all times should also be applied to the Fock–Stueckelberg theory. This means that we must require that the time derivatives of the constraints derived, for example, from the action given in Eq. (31) vanish on shell. This may lead to a fixation of multipliers $\lambda$, to redundant constraints or to new constraints. In case new constraints appear, we must again require they hold at all times.
@@ III.2 Constants of Nature as spacetime fields
$S_{\rm new}$ $=\int_{\tau_{0}}^{\tau_{1}}\!\mathrm{d}\tau\Big{\{}\mathcal{A}_{\mu}[z,\mathscr{P}]\dot{z}^{\mu}+\mathscr{P}_{i}\dot{\mathscr{Q}}^{i}$ (32)
$\ \ \ -H_{0}(\tau;z,\mathscr{P}]-\lambda^{a}\phi_{a}[z,\mathscr{P}]\Big{\}}\,,$
which is identical to Eq. (4) except for the fact that the constants of Nature $\mathscr{P}_{i}$ are promoted to fields (degrees of freedom), which are interpreted as the canonical momenta conjugate to cyclic degrees of freedom $\mathscr{Q}^{i}$. This then implies that the $\mathscr{P}$ fields do not change over time, and any initial condition will be frozen. If one assumes an initial value for the $\mathscr{P}$ fields that is constant across space (if $d>0$), then due to $\mathscr{P}$ conservation, the $\mathscr{P}$ solution to the field equations will be spatially constant at all times. On the other hand, an initial condition that changes across space would lead to “constants” of Nature with a static spatial variation. In any case, the phase space is extended to include the canonical pairs $(\mathscr{Q},\mathscr{P})$, and the new theory described by the action (32) does not exhibit constants of Nature (theory-defining parameters) but rather only fields, some of which are conserved in time. Of course, we also assume that the constraints $\phi_{a}$ obey the regularity conditions with respect to the enlarged set of fields $(z,\mathscr{P},\mathscr{Q})$.
@@ III.2.1 Parametrization
First, if $d>0$, we now define $\mathtt{Z}=(z,\ldots,\partial^{n}z,\mathscr{P},\ldots,\partial^{n}\mathscr{P})$, and we assume that the unconstrained Hamiltonian can be written as (cf. Footnote 2)
$H_{0}(\tau;z,\mathscr{P}]=\int\mathrm{d}^{d}x\ \mathscr{H}_{0}(\tau,\mathbf{x};{\mathtt{Z}(\tau,\mathbf{x})})\ ,$ (33)
@@
$\{\phi_{a}{(\mathtt{Z}(\tau,\mathbf{x}))},\mathscr{H}_{0}(\tau,\mathbf{x}^{\prime};{\mathtt{Z}(\tau,\mathbf{x}^{\prime})})\}$ (34)
$\ \ =\sum_{b=1}^{N}\int\mathrm{d}^{d}y\,\mathscr{V}_{a}^{b}(\tau,\mathbf{x},\mathbf{x}^{\prime},\mathbf{y};z,\mathscr{P}]\phi_{b}({\mathtt{Z}(\tau,\mathbf{y})})\ ,$
@@
then Eqs. (33) and (LABEL:eq:field-first-class) yield
$\{\phi_{a}({\mathtt{Z}(\tau,\mathbf{x})}),H_{0}(\tau;z,\mathscr{P}]\}$ (36)
$\ \ \ =\sum_{b=1}^{N}\int\mathrm{d}^{d}y\,\mathcal{V}_{a}^{b}(\tau,\mathbf{x},\mathbf{y};z,\mathscr{P}]\phi_{b}({\mathtt{Z}(\tau,\mathbf{y})})\ ,$
@@
$\phi_{0}=p_{T}+\mathscr{H}_{0}(T;\mathbf{x};{\mathtt{Z}})\ ,$ (37)
@@
If we include the $(T,p_{T})$ pair in the definition of the Poisson bracket in the usual way, it is straightforward to see that $\phi_{0}$ is a first-class constraint due to the original first-class algebra given by Eqs. (12) and (LABEL:eq:field-first-class), and due to the fact that the $\phi_{a}$ constraints are assumed to have no explicit dependence on time (or $T$). We can then redefine the multipliers $\lambda_{\rm new}\coloneqq(\lambda^{0},\lambda_{\rm old})$ and constraints $\phi_{\rm new}=(\phi_{0},\phi_{\rm old})$, so as to obtain a new first-class algebra with no unconstrained Hamiltonian. As a matter of notation, we can also add $p_{T}$ to the list of $\mathcal{P}$ fields, and, correspondingly $T$ to the list of $\mathcal{Q}$ fields. [^15] This is just for convenience because, in general, $T$ will not be cyclic and $p_{T}$ will not be constant in time. If $\mathscr{H}_{0}$ or $H_{0}$ depend on $T$ explicitly, we have $\dot{p}_{T}(\tau,\mathbf{x})\approx-\int\mathrm{d}^{d}y\lambda^{0}(\tau,\mathbf{y})\delta\mathscr{H}_{0}|_{\mathbf{y}}/\delta T(\tau,\mathbf{x})$ if $d>0$ or $\dot{p}_{T}(\tau)\approx-\lambda^{0}(\tau)\partial H_{0}/\partial T$ if $d=0$. In this way, the parametrized action in Eq. (39) is equivalent to
$S_{\rm new}^{\rm par}$ $=\int_{\tau_{0}}^{\tau_{1}}\!\mathrm{d}\tau\left\{\mathcal{A}_{\mu}\dot{z}^{\mu}+\mathscr{P}_{i}\dot{\mathscr{Q}}^{i}-\lambda^{a}\phi_{a}[z,\mathscr{P}]\right\},$ (40)
where the indices now run over the appropriate values for the enlarged set of variables together with implicit spatial integrations if $d>0$. We also assume that the enlarged set of constraints obeys the regularity conditions with respect to the $(z,\mathscr{P},\mathscr{Q})$ fields.
@@ III.2.2 Solving the constraints in terms of constants of Nature
$\mathscr{P}_{a}$ $\approx-\mathscr{H}_{a}(z,{\ldots,\partial^{n}z,}\mathscr{P}_{i>a}{,\ldots,\partial^{n}\mathscr{P}_{i>a}})\ ,$ (41)
$\ (a=1,\ldots,P)\ ,$
in a region of phase space. One can then use these solutions recursively to express $\mathscr{H}_{a}$ only in terms of the $z$ fields and their derivatives. In the region of phase space where this holds, the constraint hypersurface can be defined by the new set of constraints [^18] The parametrization constraint in Eq. (37) is not necessarily a particular case of Eq. (42) because $p_{T}$ need not be constant in time, i.e., a constant of Nature (cf. Footnote 15). Furthermore, notice that while the constraints that come from the parametrization procedure, such as the one in Eq. (37), may be globally valid in phase space, the solutions in Eqs. (42) may hold only in a certain region. Thus, in general, a parametrized theory with constraints of the form given in Eq. (42) may not be globally equivalent to gauge theory with constraints $\phi_{a}$, with the equivalence holding locally in certain cases. We also disregard a possible explicit spatial dependence of $\mathscr{H}_{a}$ in Eqs. (41) and (42).
$\bar{\phi}_{a}$ $\coloneqq\mathscr{P}_{a}+\mathscr{H}_{a}(z,{\ldots,\partial^{n}z})\approx 0\ ,\ (a=1,\ldots,P)\ ,$ (42)
@@ III.2.3 General constraint redefinitions
What can we conclude about the constraint transformation in Eqs. (42)? If we assume that this transformation is invertible in a region of phase space in a neighborhood of the constraint hypersurface, then it is a particular case of a general transformation $\phi_{a}\mapsto\bar{\phi}_{a}(\phi)$ of constraints that is well defined and invertible in a neighborhood of the constraint hypersurface and satisfies $\bar{\phi}_{a}=0$ if, and only if, $\phi_{a}=0$ (for every $a=1,\ldots,N$). Any such transformation must satisfy $\bar{\phi}_{a}=\tensor{M}{{}_{a}^{b}}\phi_{b}$ due to Eq. (11), with $\tensor{M}{{}_{a}^{b}}$ being the coefficients of a matrix that must be non-divergent and invertible at least in a neighborhood of the constraint hypersurface. The new constraints $\bar{\phi}_{a}$ are thus independent, and they determine the same hypersurface as the original $\phi_{a}$ functions, but they will generally have different structure functions.
@@
As a particular case, we assume that the new constraints in Eqs. (42) are functionally independent and that they obey the regularity conditions. To verify the relation $\bar{\phi}_{a}=\tensor{M}{{}_{a}^{b}}\phi_{b}$, one may expand $\phi_{a}$ ($a=1,\ldots,P$) in a Taylor series in powers of $\mathscr{P}$ to find, in particular,
$\phi_{a}=\phi_{a}|_{\bar{\phi}=0}+\frac{\partial\phi_{a}}{\partial\mathscr{P}_{b}}\bar{\phi}_{b}+\frac{1}{2}\frac{\partial^{2}\phi_{a}}{\partial\mathscr{P}_{b}\mathscr{P}_{c}}\bar{\phi}_{b}\bar{\phi}_{c}+\mathcal{O}(\bar{\phi}^{3})\,,$ (43)
where $a,b=1,\ldots,P$, $\bar{\phi}_{a}$ are given by Eq. (42), and we suppressed the arguments of $\phi_{a}$ for brevity. Then, since $\phi_{a}|_{\bar{\phi}}=0$, we have
@@
As we assume that $\phi_{a}\approx 0$ ($a=1,\ldots,P$) if, and only if, the constraints in Eq. (41) hold, then the matrix $\tensor{\left(M^{-1}\right)}{{}_{a}^{b}}$ must be invertible on the constraint hypersurface. In a neighborhood of this hypersurface, for $a,b,c=1,\ldots,P$, we can define $\frac{1}{2}\partial^{2}\phi_{a}/\partial\mathscr{P}_{b}\partial\mathscr{P}_{c}$ to be the element of row $a$ and column $b$ of a matrix $\mathtt{H}^{c}$ for each value of $c$, so that $\det M^{-1}=\det M^{-1}|_{\bar{\phi}=0}(1+\bar{\phi}_{c}\mathrm{Tr}(M\mathtt{H}^{c})|_{\bar{\phi}=0}+\mathcal{O}(\bar{\phi}^{2}))$, which is nonvanishing for $\bar{\phi}$ with sufficiently small but nonzero absolute values. We can then trivially extend $\tensor{(M^{-1})}{{}_{a}^{b}}$ to $a,b=1,\ldots,N$ because the constraints $\phi_{a}$ with $a=P+1,\ldots,N$ are not changed in this particular case. The theory can thus be defined from the action
$S_{\rm new}^{\rm depar}$ $=\int_{\tau_{0}}^{\tau_{1}}\!\mathrm{d}\tau\left\{\mathcal{A}_{\mu}\dot{z}^{\mu}+\mathscr{P}_{i}\dot{\mathscr{Q}}^{i}-\bar{\lambda}^{a}\bar{\phi}_{a}[z,\mathscr{P}]\right\},\!$ (45)